In a mixture of neon and nitrogen gases, the partial pressure of the neon is 35.5 kPa and the partial pressure of nitrogen is 23.8 kPa. There are 20.0 moles of gas present. How many moles of nitrogen are present?


1
Expert's answer
2021-04-08T12:16:40-0400

= 23.8 + 35.5 = 59.3

= 23.8/59.3 × 20

= 8.027 moles of nitrogen